# Confidence-Guided Human-AI Collaboration: Reinforcement Learning with Distributional Proxy Value Propagation for Autonomous Driving

### v1

@@ II-A Human-AI Collaboration
HAC extends the standard RL by integrating human expert into the learning process. During the training process, the human expert monitors the agent-environment interactions, and can intervene the agent’s exploration by providing guidance. Specifically, if the agent encounters a risky situation or makes a suboptimal decision, the human expert can overwrite the agent action $a_{g}$ with their own action $a_{h}$. The action applied to the environment can thus be expressed as $a_{b}=I\left(s,a\right)a_{h}+\left(1-I\left(s,a\right)\right)a_{g}$, where $I\left(s,a_{g}\right)$ is a Boolean indicator. Thus, with the human policy denoted by $\pi_{h}$, the shared policy $\pi_{b}$, which generates the actual trajectory, is defined as
$\pi_{b}(a\mid s)=\pi_{g}(a\mid s)(1-I(s,a))+\pi_{h}(a\mid s)G(s),$ (6)
@@ II-B Distributional Proxy Value Propagation
$J_{\mathcal{Z}^{g}}^{H}(\theta)=\underset{(s,a_{h},a_{n})\sim\mathcal{B}_{h}}{\mathbb{E}}\left[D_{\mathrm{KL}}\left(\delta_{1}(\cdot),\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a_{h})\right)\right],$ (12)
$J_{\mathcal{Z}^{g}}^{N}(\theta)=\underset{(s,a_{h},a_{n})\sim\mathcal{B}_{h}}{\mathbb{E}}\left[D_{\mathrm{KL}}\left(\delta_{-1}(\cdot),\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a_{g})\right)\right].$ (13)
@@
$\nabla_{\theta}J_{\mathcal{Z}^{g}}^{H}(\theta)=$ $\mathbb{E}\left[\nabla_{\theta}\frac{\left(1-Q_{\theta}(s,a)\right)^{2}}{2\sigma_{\theta}(s,a)^{2}}+\eta\frac{\nabla_{\theta}\sigma_{\theta}(s,a)}{\sigma_{\theta}(s,a)}\right]$ (14)
$=$ $\mathbb{E}[-\frac{\left(1-Q_{\theta}(s,a)\right)}{\sigma_{\theta}(s,a)^{2}}\nabla_{\theta}Q_{\theta}(s,a)$
$-\frac{\left(1-Q_{\theta}(s,a)\right)^{2}-\sigma_{\theta}(s,a)^{2}}{\sigma_{\theta}(s,a)^{3}}\eta\nabla_{\theta}\sigma_{\theta}(s,a)],$
@@
$\nabla_{\theta}J_{\mathcal{Z}^{g}}^{N}(\theta)=$ $\mathbb{E}\left[\nabla_{\theta}\frac{\left(1-Q_{\theta}(s,a)\right)^{2}}{2\sigma_{\theta}(s,a)^{2}}+\eta\frac{\nabla_{\theta}\sigma_{\theta}(s,a)}{\sigma_{\theta}(s,a)}\right]$ (15)
$=$ $\mathbb{E}\left[\frac{\left(1-Q_{\theta}(s,a)\right)}{\sigma_{\theta}(s,a)^{2}}\nabla_{\theta}Q_{\theta}(s,a)\right.$
$\left.-\frac{\left(1-Q_{\theta}(s,a)\right)^{2}-\sigma_{\theta}(s,a)^{2}}{\sigma_{\theta}(s,a)^{3}}\eta\nabla_{\theta}\sigma_{\theta}(s,a)\right],$
@@
where $\bar{\theta}$ and $\bar{\phi}$ are the target-network parameters, and $\mathcal{B}$ refers to $\mathcal{B}_{n}\cup\mathcal{B}_{h}$. Since $\widehat{\mathcal{T}}_{\mathcal{D}}^{\pi^{g}_{\bar{\Phi}}}\mathcal{Z}^{g}_{\bar{\theta}}(\cdot\mid s,a)$ is unknown, a sample-based version of (16) is applied:
$J^{TD}_{\mathcal{Z}^{g}}(\theta)=-\underset{\begin{subarray}{c}\left(s,a,s^{\prime}\right)\sim\mathcal{B},a^{\prime}\sim\pi^{g}_{\bar{\Phi}},\\ Z\left(s^{\prime},a^{\prime}\right)\sim\mathcal{Z}^{g}_{\bar{\theta}}\left(\cdot\mid s^{\prime},a^{\prime}\right)\end{subarray}}{\mathbb{E}}\left[\log\mathcal{P}\left(\widehat{y}_{z}\mid\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a)\right)\right],$ (17)
@@
$\widehat{y}_{z}=\gamma\left(Z\left(s^{\prime},a^{\prime}\right)-\alpha\log\pi^{g}_{\bar{\phi}}\left(a^{\prime}\mid s^{\prime}\right)\right).$ (18)
@@
$J_{\pi^{g}}(\phi)$ $=\underset{\begin{subarray}{c}s\sim\mathcal{B},\\ a\sim\pi^{g}_{\phi}\end{subarray}}{\mathbb{E}}\left[\underset{Z(s,a)\sim\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a)}{\mathbb{E}}[Z(s,a)]-\alpha\log\left(\pi^{g}_{\phi}(a\mid s)\right)\right]$ (21)
@@
The expected cumulative probability of failure $V_{\pi_{b}}$ of the behavior policy $\pi_{b}$ in D-PVP can be upper-bounded by:
$V_{\pi_{b}}\;\leq\;\frac{1}{1-\gamma}\Bigl{(}\epsilon+\kappa\;+\;\frac{\gamma\,\epsilon^{2}}{1-\gamma}\,K^{\prime}\Bigr{)},$ (23)
@@ II-C Shared Control Mechanism with Policy Confidence Evaluation
$\pi_{b}\left(a\mid s\right)=\pi^{r}_{\psi}\left(a\mid s\right)\left(1-\mathcal{T}_{c}(s)\right)+\pi^{g}_{\phi}\left(a\mid s\right)\mathcal{T}_{c}(s).$ (24)
@@
$J_{\pi^{r}}(\psi)$ $=\underset{s\sim\mathcal{B},a\sim\pi_{\psi}^{r}}{\mathbb{E}}\bigg{[}\underset{Z(s,a)\sim\mathcal{Z}^{c}_{\zeta}(\cdot\mid s,a)}{\mathbb{E}}[Z(s,a)]$ (25)
@@
$J^{TD}_{\mathcal{Z}^{c}}(\zeta)=-\underset{\begin{subarray}{c}\left(s,a,r,s^{\prime}\right)\sim\mathcal{B},a^{\prime}\sim\pi^{r}_{\bar{\psi}},\\ Z\left(s^{\prime},a^{\prime}\right)\sim\mathcal{Z}^{c}_{\bar{\zeta}}\left(\cdot\mid s^{\prime},a^{\prime}\right)\end{subarray}}{\mathbb{E}}\left[\log\mathcal{P}\left(y_{z}^{r}\mid\mathcal{Z}^{c}_{\zeta}(\cdot\mid s,a)\right)\right],$ (26)
@@
$y_{z}^{r}=r+\gamma\left(Z\left(s^{\prime},a^{\prime}\right)-\alpha\log\pi^{r}_{\bar{\psi}}\left(a^{\prime}\mid s^{\prime}\right)\right).$ (27)
@@
(Bounded variance of return distributions.) To simplify the analysis, the variances of the return distributions for both the human-guided policy ($\pi_{g}$) and the reward-guided policy ($\pi_{r}$) are bounded by a constant $\sigma_{\max}$. Specifically:
@@
With the confidence-based intervention function $\mathcal{T}_{\text{c}}(s)$ defined in (30), the return of the behavior policy $\pi_{b}$ is guaranteed to be no worse than:
$J\left(\pi_{b}\right)\geq J\left(\pi_{g}\right)-(1-\beta)\cdot\frac{\sqrt{2}\cdot\sigma_{\max}\cdot\Phi^{-1}(\delta)}{1-\gamma},$ (31)
@@
For behavior policy $\pi_{b}$ deduced by human-guided policy $\pi_{g}$, reward-guided policy $\pi_{r}$ and an intervention function $\mathcal{T}_{c}(s)$, the state distribution discrepancy between $\pi_{b}$ and $\pi_{r}$ is bounded by (Theorem 3.2 in [46])
$\left\|d_{\pi_{b}}-d_{\pi_{r}}\right\|_{1}\leq\frac{\beta\gamma}{1-\gamma}\mathbb{E}_{s\sim d_{\pi_{b}}}\left\|\pi_{g}(\cdot\mid s)-\pi_{r}(\cdot\mid s)\right\|_{1},$ (32)
where $\beta=\frac{\mathbb{E}_{s\sim d_{\pi_{b}}}\left[\mathcal{T}(s)\left\|\pi_{g}(\cdot\mid s)-\pi_{r}(\cdot\mid s)\right\|_{1}\right]}{\mathbb{E}_{s\sim d_{\pi_{b}}}\left\|\pi_{g}(\cdot\mid s)-\pi_{r}(\cdot\mid s)\right\|_{1}}\in[0,1]$ is the expected intervention rate weighted by the policy discrepancy.
@@
$J\left(\pi_{b}\right)-J\left(\pi_{g}\right)$ (34)
$=\underset{s_{n},a_{n}\sim\tau_{\pi_{b}}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}A_{g}\left(s_{n},a_{n}\right)\right]$
$=\underset{s_{n},a_{n}\sim\tau_{\pi_{b}}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}\left[Q_{g}\left(s_{n},a_{n}\right)-V_{g}\left(s_{n}\right)\right]\right]$
$=\underset{s_{n}\sim\tau_{\pi_{b}}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}\left[\mathbb{E}_{a\sim\pi_{b}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)-V_{g}\left(s_{n}\right)\right]\right]$
$=\underset{s_{n}\sim\tau_{\pi_{b}}}{\mathbb{E}}\bigg{[}\sum_{n=0}^{\infty}\gamma^{n}\left[\mathcal{T}_{c}\left(s_{n}\right)\mathbb{E}_{a\sim\pi_{g}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)\right.$
$\quad\quad\quad\left.+\left(1-\mathcal{T}_{c}\left(s_{n}\right)\right)\mathbb{E}_{a\sim\pi_{r}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)-V_{g}\left(s_{n}\right)\right]\bigg{]}$
$=\underset{s_{n}\sim\tau_{\pi_{b}}}{\mathbb{E}}\bigg{[}\sum_{n=0}^{\infty}\gamma^{n}\left[\left(1-\mathcal{T}_{c}\left(s_{n}\right)\right)\left[\mathbb{E}_{a\sim\pi_{r}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-V_{g}\left(s_{n}\right)\right]\right]\bigg{]}$
$=(1-\beta)\underset{s_{n}\sim\tau_{\pi_{b}}}{\mathbb{E}}\bigg{[}\sum_{n=0}^{\infty}\gamma^{n}\left[\mathbb{E}_{a\sim\pi_{r}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-V_{g}\left(s_{n}\right)\right]\bigg{]}.$
Note that the choice of $\pi_{r}$ is guided by the confidence probability $P\left(Q_{g}\left(s,a_{r}\right)>Q_{g}\left(s,a_{g}\right)\right)$. When the human-guided policy is selected, its expected advantage over the human-guided policy can be expressed as:
$\mathbb{E}_{a\sim\pi_{r}\left(\cdot\mid s_{n}\right)}$ $Q_{g}\left(s_{n},a\right)-V_{g}\left(s_{n}\right)$ (35)
$=\mathbb{E}_{a\sim\pi_{r}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)-\mathbb{E}_{a\sim\pi_{g}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right).$
@@
$P\left(Q_{g}\left(s,a_{r}\right)>Q_{g}\left(s,a_{g}\right)\right)\leq 1-\delta,$
@@
$\mathbb{E}_{a\sim\pi_{r}\left(\cdot\mid s_{n}\right)}Q_{g}\left(s_{n},a\right)-\mathbb{E}_{a\sim\pi_{g}\left(\cdot\mid s_{n}\right)}$ $Q_{g}\left(s_{n},a\right)$ (36)
@@
where $\Phi^{-1}(\delta)$ is the inverse CDF of the standard normal distribution, and $\sigma_{r}^{2},\sigma_{g}^{2}$ are the variances of the value-distribution for $\pi_{r}$ and $\pi_{g}$, respectively.
@@
$J\left(\pi_{b}\right)-$ $J\left(\pi_{g}\right)$ (37)
$\geq-(1-\beta)\mathbb{E}_{s_{n}\sim\tau_{\pi_{b}}}\left[\sum_{n=0}^{\infty}\gamma^{n}\cdot\sqrt{\sigma_{r}^{2}+\sigma_{h}^{2}}\cdot\Phi^{-1}(\delta)\right].$
@@
$J\left(\pi_{b}\right)-J\left(\pi_{g}\right)\geq-(1-\beta)\cdot\frac{\sqrt{2}\cdot\sigma_{\max}\cdot\Phi^{-1}(\delta)}{1-\gamma}.$
@@
$J\left(\pi_{b}\right)\geq J\left(\pi_{g}\right)-(1-\beta)\cdot\frac{\sqrt{2}\cdot\sigma_{\max}\cdot\Phi^{-1}(\delta)}{1-\gamma}.$

### v2

@@ II-A Human-AI Collaboration
HAC extends the standard RL by integrating human expert into the learning process. During the training process, the human expert monitors the agent-environment interactions, and can intervene the agent’s exploration by providing guidance. Specifically, if the agent encounters a risky situation or makes a suboptimal decision, the human expert can overwrite the agent action $a_{g}$ with their own action $a_{h}$. The action applied to the environment can thus be expressed as $\widehat{a}=I\left(s,a\right)a_{h}+\left(1-I\left(s,a\right)\right)a_{g}$, where $I\left(s,a_{g}\right)$ is a Boolean indicator. Thus, with the human policy denoted by $\pi^{h}$, the shared policy $\pi^{b}$, which generates the actual trajectory, is defined as
$\pi^{b}(a\mid s)=\pi^{g}(a\mid s)(1-I(s,a))+\pi^{h}(a\mid s)G(s),$ (6)
@@ II-B Distributional Proxy Value Propagation
$J_{\mathcal{Z}^{g}}^{H}(\theta)=\underset{(s,a_{h},a_{g})\sim\mathcal{B}_{h}}{\mathbb{E}}\left[D_{\mathrm{KL}}\left(\delta_{1}(\cdot),\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a_{h})\right)\right],$ (12)
$J_{\mathcal{Z}^{g}}^{N}(\theta)=\underset{(s,a_{h},a_{g})\sim\mathcal{B}_{h}}{\mathbb{E}}\left[D_{\mathrm{KL}}\left(\delta_{-1}(\cdot),\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a_{g})\right)\right].$ (13)
@@
$\nabla_{\theta}J_{\mathcal{Z}^{g}}^{H}(\theta)=$ $\mathbb{E}\left[\nabla_{\theta}\frac{\left(1-Q_{\theta}(s,a_{h})\right)^{2}}{2\sigma_{\theta}(s,a_{h})^{2}}+\eta\frac{\nabla_{\theta}\sigma_{\theta}(s,a_{h})}{\sigma_{\theta}(s,a_{h})}\right]$ (14)
$=$ $\mathbb{E}[-\frac{\left(1-Q_{\theta}(s,a_{h})\right)}{\sigma_{\theta}(s,a_{h})^{2}}\nabla_{\theta}Q_{\theta}(s,a_{h})$
$-\frac{\left(1-Q_{\theta}(s,a_{h})\right)^{2}-\sigma_{\theta}(s,a_{h})^{2}}{\sigma_{\theta}(s,a_{h})^{3}}\eta\nabla_{\theta}\sigma_{\theta}(s,a_{h})],$
@@
$\nabla_{\theta}J_{\mathcal{Z}^{g}}^{N}(\theta)=$ $\mathbb{E}\left[\nabla_{\theta}\frac{\left(-1-Q_{\theta}(s,a_{g})\right)^{2}}{2\sigma_{\theta}(s,a_{g})^{2}}+\eta\frac{\nabla_{\theta}\sigma_{\theta}(s,a_{g})}{\sigma_{\theta}(s,a_{g})}\right]$ (15)
$=$ $\mathbb{E}\left[\frac{\left(1+Q_{\theta}(s,a_{g})\right)}{\sigma_{\theta}(s,a_{g})^{2}}\nabla_{\theta}Q_{\theta}(s,a_{g})\right.$
$\left.-\frac{\left(1+Q_{\theta}(s,a_{g})\right)^{2}-\sigma_{\theta}(s,a_{g})^{2}}{\sigma_{\theta}(s,a_{g})^{3}}\eta\nabla_{\theta}\sigma_{\theta}(s,a_{g})\right],$
@@
where $\bar{\theta}$ and $\bar{\phi}$ are the target-network parameters, and $\mathcal{B}$ refers to $\mathcal{B}_{g}\cup\mathcal{B}_{h}$. Since $\widehat{\mathcal{T}}_{\mathcal{D}}^{\pi^{g}_{\bar{\Phi}}}\mathcal{Z}^{g}_{\bar{\theta}}(\cdot\mid s,a)$ is unknown, a sample-based version of (16) is applied:
$J^{TD}_{\mathcal{Z}^{g}}(\theta)=-\underset{\begin{subarray}{c}\left(s,a,s^{\prime}\right)\sim\mathcal{B},a^{\prime}\sim\pi^{g}_{\bar{\Phi}},\\ Z^{g}\left(s^{\prime},a^{\prime}\right)\sim\mathcal{Z}^{g}_{\bar{\theta}}\left(\cdot\mid s^{\prime},a^{\prime}\right)\end{subarray}}{\mathbb{E}}\left[\log\mathcal{P}\left(\widehat{y}_{z}\mid\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a)\right)\right],$ (17)
@@
$\widehat{y}_{z}=\gamma\left(Z^{g}\left(s^{\prime},a^{\prime}\right)-\alpha\log\pi^{g}_{\bar{\phi}}\left(a^{\prime}\mid s^{\prime}\right)\right).$ (18)
@@
$J_{\pi^{g}}(\phi)$ $=\underset{\begin{subarray}{c}s\sim\mathcal{B},\\ a\sim\pi^{g}_{\phi}\end{subarray}}{\mathbb{E}}\left[\underset{Z^{g}(s,a)\sim\mathcal{Z}^{g}_{\theta}(\cdot\mid s,a)}{\mathbb{E}}[Z^{g}(s,a)]-\alpha\log\left(\pi^{g}_{\phi}(a\mid s)\right)\right]$ (21)
@@
The expected cumulative probability of failure $V_{\pi^{b}}$ of the behavior policy $\pi^{b}$ in D-PVP can be upper-bounded by:
$V_{\pi^{b}}\;\leq\;\frac{1}{1-\gamma}\Bigl{(}\epsilon+\kappa\;+\;\frac{\gamma\,\epsilon^{2}}{1-\gamma}\,K^{\prime}\Bigr{)},$ (23)
@@ II-C Shared Control Mechanism with Policy Confidence Evaluation
$\pi^{b}\left(a\mid s\right)=\pi^{r}_{\psi}\left(a\mid s\right)\left(1-\mathcal{T}_{c}(s)\right)+\pi^{g}_{\phi}\left(a\mid s\right)\mathcal{T}_{c}(s).$ (24)
@@
$J_{\pi^{r}}(\psi)$ $=\underset{s\sim\mathcal{B},a\sim\pi_{\psi}^{r}}{\mathbb{E}}\bigg{[}\underset{Z^{c}(s,a)\sim\mathcal{Z}^{c}_{\zeta}(\cdot\mid s,a)}{\mathbb{E}}[Z^{c}(s,a)]$ (25)
@@
$J^{TD}_{\mathcal{Z}^{c}}(\zeta)=-\underset{\begin{subarray}{c}\left(s,a,r,s^{\prime}\right)\sim\mathcal{B},a^{\prime}\sim\pi^{r}_{\bar{\psi}},\\ Z^{c}\left(s^{\prime},a^{\prime}\right)\sim\mathcal{Z}^{c}_{\bar{\zeta}}\left(\cdot\mid s^{\prime},a^{\prime}\right)\end{subarray}}{\mathbb{E}}\left[\log\mathcal{P}\left(y_{z}^{r}\mid\mathcal{Z}^{c}_{\zeta}(\cdot\mid s,a)\right)\right],$ (26)
@@
$y_{z}^{r}=r+\gamma\left(Z^{c}\left(s^{\prime},a^{\prime}\right)-\alpha\log\pi^{r}_{\bar{\psi}}\left(a^{\prime}\mid s^{\prime}\right)\right).$ (27)
@@
(Bounded variance of return distributions.) To simplify the analysis, the variances of the return distributions for both the human-guided policy ($\pi^{g}$) and the reward-guided policy ($\pi^{r}$) are bounded by a constant $\sigma_{\max}$. Specifically:
@@
With the confidence-based intervention function $\mathcal{T}_{\text{c}}(s)$ defined in (30), the return of the behavior policy $\pi^{b}$ is guaranteed to be no worse than:
$J\left(\pi^{b}\right)\geq J\left(\pi^{g}\right)-(1-\beta)\cdot\frac{\sqrt{2}\cdot\sigma_{\max}\cdot\Phi^{-1}(\delta)}{1-\gamma},$ (31)
@@
For behavior policy $\pi^{b}$ deduced by human-guided policy $\pi^{g}$, reward-guided policy $\pi^{r}$ and an intervention function $\mathcal{T}_{c}(s)$, the state distribution discrepancy between $\pi^{b}$ and $\pi^{r}$ is bounded by (Theorem 3.2 in [46])
$\left\|d_{\pi^{b}}-d_{\pi^{r}}\right\|_{1}\leq\frac{\beta\gamma}{1-\gamma}\mathbb{E}_{s\sim d_{\pi^{b}}}\left\|\pi^{g}(\cdot\mid s)-\pi^{r}(\cdot\mid s)\right\|_{1},$ (32)
where $\beta=\frac{\mathbb{E}_{s\sim d_{\pi^{b}}}\left[\mathcal{T}(s)\left\|\pi^{g}(\cdot\mid s)-\pi^{r}(\cdot\mid s)\right\|_{1}\right]}{\mathbb{E}_{s\sim d_{\pi^{b}}}\left\|\pi^{g}(\cdot\mid s)-\pi^{r}(\cdot\mid s)\right\|_{1}}\in[0,1]$ is the expected intervention rate weighted by the policy discrepancy.
@@
$J\left(\pi^{b}\right)-J\left(\pi^{g}\right)$ (34)
$=\underset{s_{n},a_{n}\sim\tau_{\pi^{b}}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}A^{g}\left(s_{n},a_{n}\right)\right]$
$=\underset{s_{n},a_{n}\sim\tau_{\pi^{b}}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}\left[Q^{g}\left(s_{n},a_{n}\right)-V^{g}\left(s_{n}\right)\right]\right]$
$=\underset{s_{n}\sim\tau_{\pi^{b}}}{\mathbb{E}}\left[\sum_{n=0}^{\infty}\gamma^{n}\left[\mathbb{E}_{a\sim\pi^{b}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)-V^{g}\left(s_{n}\right)\right]\right]$
$=\underset{s_{n}\sim\tau_{\pi^{b}}}{\mathbb{E}}\bigg{[}\sum_{n=0}^{\infty}\gamma^{n}\left[\mathcal{T}_{c}\left(s_{n}\right)\mathbb{E}_{a\sim\pi^{g}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)\right.$
$\quad\quad\quad\left.+\left(1-\mathcal{T}_{c}\left(s_{n}\right)\right)\mathbb{E}_{a\sim\pi^{r}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)-V^{g}\left(s_{n}\right)\right]\bigg{]}$
$=\underset{s_{n}\sim\tau_{\pi^{b}}}{\mathbb{E}}\bigg{[}\sum_{n=0}^{\infty}\gamma^{n}\left[\left(1-\mathcal{T}_{c}\left(s_{n}\right)\right)\left[\mathbb{E}_{a\sim\pi^{r}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.-V^{g}\left(s_{n}\right)\right]\right]\bigg{]}$
$=(1-\beta)\underset{s_{n}\sim\tau_{\pi^{b}}}{\mathbb{E}}\bigg{[}\sum_{n=0}^{\infty}\gamma^{n}\left[\mathbb{E}_{a\sim\pi^{r}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-V^{g}\left(s_{n}\right)\right]\bigg{]}.$
Note that the choice of $\pi^{r}$ is guided by the confidence probability $P\left(Q^{g}\left(s,a_{r}\right)>Q^{g}\left(s,a_{g}\right)\right)$. When the human-guided policy is selected, its expected advantage over the human-guided policy can be expressed as:
$\mathbb{E}_{a\sim\pi^{r}\left(\cdot\mid s_{n}\right)}$ $Q^{g}\left(s_{n},a\right)-V^{g}\left(s_{n}\right)$ (35)
$=\mathbb{E}_{a\sim\pi^{r}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)-\mathbb{E}_{a\sim\pi^{g}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right).$
@@
$P\left(Q^{g}\left(s,a_{r}\right)>Q^{g}\left(s,a_{g}\right)\right)\leq 1-\delta,$
@@
$\mathbb{E}_{a\sim\pi^{r}\left(\cdot\mid s_{n}\right)}Q^{g}\left(s_{n},a\right)-\mathbb{E}_{a\sim\pi^{g}\left(\cdot\mid s_{n}\right)}$ $Q^{g}\left(s_{n},a\right)$ (36)
@@
where $\Phi^{-1}(\delta)$ is the inverse CDF of the standard normal distribution, and $\sigma_{r}^{2},\sigma_{g}^{2}$ are the variances of the value-distribution for $\pi^{r}$ and $\pi^{g}$, respectively.
@@
$J\left(\pi^{b}\right)-$ $J\left(\pi^{g}\right)$ (37)
$\geq-(1-\beta)\mathbb{E}_{s_{n}\sim\tau_{\pi^{b}}}\left[\sum_{n=0}^{\infty}\gamma^{n}\cdot\sqrt{\sigma_{r}^{2}+\sigma_{h}^{2}}\cdot\Phi^{-1}(\delta)\right].$
@@
$J\left(\pi^{b}\right)-J\left(\pi^{g}\right)\geq-(1-\beta)\cdot\frac{\sqrt{2}\cdot\sigma_{\max}\cdot\Phi^{-1}(\delta)}{1-\gamma}.$
@@
$J\left(\pi^{b}\right)\geq J\left(\pi^{g}\right)-(1-\beta)\cdot\frac{\sqrt{2}\cdot\sigma_{\max}\cdot\Phi^{-1}(\delta)}{1-\gamma}.$